A bottle contains 750 mL of fruit punch with a concentration of 50% pure fruit juice. Jill drinks 100 mL of the punch and then refills the bottle with an equal amount of a cheaper brand of punch. If the concentration of juice in the bottle is now reduced to 48%, what was the concentration in the punch that Jill added?
Answer by stanbon(75887)

A bottle contains 750 mL of fruit punch with a concentration of 50% pure fruit juice. Jill drinks 100 mL of the punch and then refills the bottle with an equal amount of a cheaper brand of punch. If the concentration of juice in the bottle is now reduced to 48%, what was the concentration in the punch that Jill added?
-------
fruit + fruit = fruit
0.50*650 + x*100 = 0.48(650+100)
------
50*650 + 10000x = 48*650 + 48*100
-------
10000x = 3500
----
x = 0.35 = 35% (concentration of added punch)
=========================
Cheers,
Stan H.
==================